Grade 12General Physics

A water fountain on the ground sprinkles water all around it. If the speed of water coming out of the fountain is v, the total area around the fountain that gets wet is

sudhanshu
12 Years agoGrade 12

3 Answers

Joshua
9 Years ago
Maximum range, R=v^2g. = Radius of circleTotal area around the fountain that gets wet,A=πR^2 =π[v^2/g]^2 =πv^4/g^2
Raza
7 Years ago
We know that for maximum range 
      R = v^2/g = area of circle
   And the area that gets wet by fountain is 
      A = πR^2
    Putting the values of R; we get;
      A = π * (v^2/g)^2
      Answer is : A = π * v^4/g^2.
     
Rishi Sharma
5 Years ago
Dear Student,
Please find below the solution to your problem.

We know that for maximum range
R = v^2/g = area of circle
And the area that gets wet by fountain is
A = πR^2
Putting the values of R; we get;
A = π * (v^2/g)^2
Answer is : A = π * v^4/g^2

Thanks and Regards